What is the molar concentration of a 2% by mass H2SO4 solution with density of 1.24gmL?


1
Expert's answer
2021-03-17T06:40:54-0400

Molar Mass of H2SO4= (2×1)+(32×1)+(16×4)

=98g/mol

C=n/v

V= mtot/p

= 0.02×1.24gml/98g/mol

= 2.53×10^-4 mol/L or 0.000253mol/L